OpenStudy (anonymous):

how do you do this???? (√1/√7 ) (√7/√11)

OpenStudy (anonymous):

is this multiplication

OpenStudy (anonymous):

yes, but how do you do it????? they are fractions with square root signs!

OpenStudy (helder_edwin):

\[ \large \frac{\sqrt{1}}{\sqrt{7}}\cdot\frac{\sqrt{7}}{\sqrt{11}} \] u can cancel the \(\sqrt{7}\)'s and \(\sqrt{1}=1\)

OpenStudy (anonymous):

so it would be 1/ √11? and then i would simplify?

OpenStudy (helder_edwin):

no that's it. unless u have to rationalize the expresion.

OpenStudy (radar):

 If rationalizing is required.